### v1

# CAN-STRESS: A Real-World Multimodal Dataset for Understanding Cannabis Use, Stress, and Physiological Responses

Abstract
Coping with stress is one of the most frequently cited reasons for chronic cannabis use. Therefore, it is hypothesized that cannabis users exhibit distinct physiological stress responses compared to non-users, and these differences would be more pronounced during moments of consumption. However, there is a scarcity of publicly available datasets that allow such hypotheses to be tested in real-world environments. This paper introduces a dataset named CAN-STRESS, collected using Empatica E4 wristbands. The dataset includes physiological measurements such as skin conductance, heart rate, and skin temperature from 82 participants (39 cannabis users and 43 non-users) as they went about their daily lives. Additionally, the dataset includes self-reported surveys where participants documented moments of cannabis consumption, exercise, and rated their perceived stress levels during those moments. In this paper, we publicly release the CAN-STRESS dataset, which we believe serves as a highly reliable resource for examining the impact of cannabis on stress and its associated physiological markers.
I Introduction
Due to its widespread consumption and the increasing legalization and decriminalization in many regions, studying cannabis use and its implications is essential [1]. Understanding the physiological and psychological effects of cannabis is crucial for informing policy, healthcare, and individual decision-making. Furthermore, differentiating between the physiological responses of cannabis users and non-users offers insights into how chronic use may alter stress regulation, a frequently cited reason for cannabis consumption [2]. However, most existing studies are conducted in controlled laboratory environments, where ecological validity is constrained [3]. These settings fail to capture the nuanced, real-world conditions under which individuals consume cannabis, including the interplay of stressors and other contextual factors. Bridging this gap through field-based research is imperative to develop a comprehensive understanding of the effects of cannabis’ under realistic settings and enhance the applicability of findings to broader populations and real-world scenarios.
Figure 1: The process of collecting physiological data using wearable devices, storing it in a centralized web portal, and enabling access for various research applications addresses the limitations of laboratory-based studies. The stored data can be used to advance research in fields such as stress detection, activity recognition, and medical diagnostics.
Recently, wearable devices have shown great potential for studying individuals as they go about their daily routines[4, 5]. Many can collect various data modalities simultaneously as they are equipped with multiple sensors [6, 7]. Their small size and user-friendly design minimize disruption to participants’ everyday activities, thereby reducing the introduction of confounding variables during subsequent analyses[8, 9, 10]. Moreover, their wireless connectivity provides researchers with real-time control, which closely approximates the level of oversight achievable in laboratory settings[11, 12]. Additionally, their built-in processing units can handle advanced signal processing algorithms [11, 13, 14, 15].
To advance research on cannabis consumption and its physiological effects, this paper presents CAN-STRESS, a multimodal dataset gathered using wearable wristbands from 82 participants during a full day as they engaged in their usual daily activities. The data collection process, including participant recruitment, study protocols, and the use of self-reported questionnaires to document daily activities such as cannabis use, sleep, exercise, and perceived stress levels, has been thoroughly detailed in [16].
This dataset is one of the largest publicly available resources capturing the physiological characteristics of cannabis users in real-world settings. By including both cannabis users and non-users, it facilitates balanced comparisons and robust analysis. CAN-STRESS offers a range of modalities, including Electrodermal Activity (EDA), Heart Rate (HR), body temperature, and accelerometer readings, with synchronized and high-quality data that address the limitations of previous laboratory-based studies. The dataset is highly suitable for diverse research applications, such as exploring stress biomarkers, analyzing the physiological effects of cannabis consumption [16], and developing advanced signal processing algorithms[17].
In the next section, we briefly discuss the process of collecting the data. We then outline the various components of the dataset. In section four, we highlight key characteristics observed within the dataset. We then present some applications and preliminary results derived from our dataset in section five. Finally, the conclusion is provided in section six.
II Data Collection
Participants for this study were recruited from the Palouse region of Washington State over a period spanning February 2022 to November 2023. To qualify, individuals needed to be at least 21 years old, proficient in English, and in possession of a smartphone. Individuals were excluded if they had a history of neurological conditions, such as head injuries or seizures, or severe psychiatric disorders like psychosis, autism, or bipolar I disorder. Furthermore, participants who reported heavy alcohol use—defined as consuming more than four drinks on more than four occasions per week—were excluded, as were those who had used illicit drugs, nicotine, or tobacco products in the past three months or were currently taking corticosteroid medications. Cannabis users in the study were required to report daily or near-daily consumption (at least four times per week) over the past year, while non-users had to have consumed cannabis fewer than ten times in their lifetime, with no use in the previous year.
Recruitment was carried out using a combination of online advertisements and physical postings in local businesses, including cannabis dispensaries, coffee shops, and laundromats. These advertisements featured a URL and a scannable QR code, both of which directed interested individuals to a brief online screening survey. At the end of the survey, eligible participants were invited to provide their email addresses if they wished to be contacted for the study. Once eligibility was confirmed by a researcher, participants were contacted to schedule an initial Zoom meeting. During this meeting, the researcher explained the study’s procedures in detail and obtained informed consent. Following the meeting, study materials were delivered to the participants’ homes, including the Empatica E4 [12], a medical-grade wearable device (figure 2), along with detailed written instructions. Participants were asked to wear the wristband on their non-dominant wrist continuously for a 24-hour period, beginning as soon as they woke up on the designated day of data collection and ending the following morning, and they were instructed not to use alcohol on the study day to avoid confounds. For additional details regarding the collection of CAN-STRESS, please refer to [16].
Figure 2: Empatica E4 wristband
III Data Modalities
Figure 3: Comparison of Dataset Features Between Users and Non-Users
Our dataset consists of two major modalities: a self-reported questionnaire and multimodal physiological data collected through the E4 wearable wristband. Together, these modalities provide a comprehensive view of participants’ activities, physiological responses, and subjective experiences. Aligning and integrating the two sources of the dataset allows researchers to identify and analyze the correlations between different events, e.g. cannabis consumption vs exercise.
III-A Self-Reported Questionnaire
The first modality consists of a structured questionnaire that participants completed during the data collection period. These self-reported entries provide timestamps and labels that can be matched with the second modality to facilitate the analysis of relationships between physiological signals and daily activities. The questionnaire captures significant moments throughout the participants’ day, including:
  • Sleep Patterns: Participants recorded the times they went to bed and woke up, which provided insight into their sleep duration and routines.
  • Cannabis Use: Participants documented the start and end times of their cannabis consumption.
  • Exercise Activities: The peak moments of physical activity during the day were also recorded, which in turn enabled us to account for physiological changes associated with these activities.
  • Stress Ratings: At these moments (e.g., cannabis use, exercise, sleep), participants rated their perceived stress on a scale of 1 to 10, which provided subjective insights into their mental state throughout the day.
III-B Physiological Data from the Wristband
The second modality includes continuous physiological data collected using the Empatica E4 wristband, a medical-grade wearable device. This modality consists of the following signals:
  • Accelerometer (ACC): Captures triaxial movement data at 32 Hz, which can be used to detect activity patterns, such as exercise or sedentary behavior.
  • Blood Volume Pulse (BVP): Measured at 64 Hz, BVP provides raw data for calculating heart rate and interbeat intervals that offers insights into cardiovascular dynamics [18].
  • Electrodermal Activity (EDA): Collected at 4 Hz, EDA measures skin conductance, which is closely associated with stress and emotional arousal [19].
  • Heart Rate (HR): Derived from BVP, HR is sampled at 1 Hz and reflects real-time changes in cardiovascular activity.
  • Interbeat Interval (IBI): Derived from BVP, IBI represents the time between successive heartbeats and is critical for heart rate variability analysis [20].
  • Body Temperature (TEMP): Recorded at 4 Hz, TEMP tracks changes in skin temperature, which may be indicative of physiological or environmental changes.
These physiological signals provide high-resolution, multivariate data that capture participants’ physical states and responses in real time. By combining the self-reported data with wristband signals, researchers can investigate the interplay between subjective experiences (e.g., stress ratings) and objective physiological responses (e.g., changes in HR or EDA) across various activities and contexts. [21]
IV Data Features
To provide a comprehensive overview of the dataset, we present key statistical features that highlight differences between cannabis users and non-users, as well as physiological variations associated with self-reported stress levels. These analyses help characterize both baseline physiological patterns and stress-related responses.
IV-A Group-Level Comparisons: Users vs. Non-Users
This subsection presents key statistical features comparing cannabis users and non-users. These features include the average recording duration, self-reported stress ratings, sleep times, and physiological signals such as EDA and HR. Additionally, we include EDA peaks per minute (PPM), a well-established metric closely associated with stress, which provides valuable insights into stress-related physiological responses. These statistics offer insights into the typical physiological and behavioral differences between the two groups and serve as a foundation for further analyses.
Table I provides a detailed comparison of cannabis users and non-users across various dimensions. For each group, we report the number of participants, the average recording time (hours), the average self-reported stress rating (1–10 scale), the average sleep duration (hours/day), the mean EDA value (microsiemens), the mean EDA PPM, and the mean HR (beats per minute). Standard deviations are also included to illustrate the variability within each group. These features are further visualized in Figure 3, which presents boxplots for each metric. This figure highlights key trends such as higher stress levels, EDA values, and PPM in users, alongside longer sleep durations in non-users.
TABLE I: Summary of Dataset Features for Users and Non-Users
Feature Users Non-Users
Number of Subjects 39 43
Average Recording Time (hrs) $22.09\pm 4.54$ $21.40\pm 5.45$
Average Stress Rating (1–10) $5.27\pm 2.48$ $4.90\pm 2.47$
Average Time Slept (hrs/day) $7.97\pm 0.83$ $8.17\pm 1.08$
Average EDA Value ($\mu$S) $1.54\pm 1.34$ $1.18\pm 1.03$
Average EDA Peaks (PPM) $1.03\pm 0.34$ $0.97\pm 0.31$
Average Heart Rate (BPM) $91.68\pm 7.52$ $88.77\pm 5.42$
IV-B Physiological Features Around Stress Events
In addition to examining differences between cannabis users and non-users, we analyzed how physiological signals change in response to self-reported stress events. Specifically, we investigated fluctuations in EDA and HR surrounding these moments. To systematically assess these changes, we stratified self-reported stress ratings into four categories: no stress (0–1), low stress (2–4), medium stress (5–7), and high stress (8–9).
For each stress-labeled moment in the dataset, we extracted 60-second segments of EDA and HR data recorded during that period and calculated the average and the standard deviation of the signals to characterize their behavior. As expected, both EDA and HR values tend to be lower in moments associated with lower reported stress levels, while higher stress levels correspond to an increase in physiological activation. Figures 4 and 5 and table II visualize these differences, showing the distribution of EDA and HR across the four stress categories. As you can see, both these signals fluctuate in a much higher range for hight rates of stress. These findings support the expected relationship between physiological arousal and stress intensity and provide a basis for further modeling and predictive analysis.
TABLE II: Mean and Standard Deviation of Physiological Signals Across Stress Levels
Signal No Stress Low Stress Mild Stress High Stress
EDA $0.85\pm 0.11$ $0.73\pm 0.06$ $1.12\pm 0.07$ $1.79\pm 0.08$
HR $88.36\pm 3.89$ $85.23\pm 2.79$ $87.58\pm 2.31$ $89.55\pm 4.63$
Figure 4: Averaged EDA signals recorded over 60-second intervals surrounding moments rated as no, low, mild, and high stress. As shown, EDA tends to be higher during moments of greater reported stress. The shaded regions around the signal represent the standard deviation which illustrates variability across subjects.
Figure 5: Averaged HR signals recorded over 60-second intervals surrounding moments rated as no, low, mild, and high stress. HR increases during high-stress moments, with a more rapid rise compared to lower stress levels. The shaded regions around the signal indicate the std of the signal, reflecting participant variability in cardiovascular responses.
V Application and Use Cases
The CAN-STRESS dataset has proven to be an invaluable resource for advancing research in physiological monitoring, cannabis consumption detection, and stress-related behaviors in real-world settings. Below, we discuss three studies that utilized this dataset and highlight how its multimodal design and real-world applicability contributed to their findings.
The first study [10] leveraged the CAN-STRESS dataset to develop a system for identifying cannabis consumption sessions in naturalistic environments. The authors of [10] focused on using Electrodermal Activity (EDA), a key physiological marker of stress, to model cannabis use detection as a supervised learning problem. By training a neural network on data collected from wearable sensors, the authors demonstrated the feasibility of detecting cannabis use with high accuracy in uncontrolled settings. Additionally, the dataset’s participant diversity and real-world conditions allowed the team to apply transfer learning techniques, which significantly improved the system’s performance for individual users. This dataset enabled the development of a robust and scalable detection framework that would have been impossible to design using laboratory-collected data alone, given its limited variability and constrained scenarios.
In another study [22], the CAN-STRESS dataset was utilized to explore the potential of self-supervised learning for detecting cannabis consumption with limited labeled data. The researchers introduced the CUDLE (Cannabis Use Detection with Label Efficiency) framework, which first used data augmentation techniques and a contrastive learning approach to train models on unlabeled wearable sensor data. The dataset’s extensive physiological measurements and user-reported moments of consumption provided the ideal foundation for validating this novel approach. By combining real-world data with self-supervised pretext tasks, the authors showcased the effectiveness of the dataset in enabling high-performance detection models while significantly reducing reliance on large-scale labeled datasets. This research highlights how CAN-STRESS supports innovation in computational health by addressing challenges unique to free-living environments.
Finally, the CAN-STRESS dataset was instrumental in a study investigating disruptions in diurnal stress rhythms among cannabis users[16]. The wearable device’s ability to capture continuous physiological signals, such as EDA and heart rate variability, alongside self-reported stress ratings, allowed researchers to assess stress-related biomarkers in real-world contexts. These naturalistic conditions were critical in revealing patterns like blunted cortisol awakening responses and elevated afternoon cortisol levels in cannabis users, which traditional lab-based studies often fail to capture. A key finding of [16] is that cortisol levels decreased after cannabis use in people’s natural environments as lab studies have found that cortisol increases after cannabis use in the lab. This finding is more consistent with the fact that many cannabis users use cannabis to decrease stress. The dataset’s inclusion of both users and non-users provided a balanced framework for comparison and enabled a more comprehensive understanding of how chronic cannabis use may affect stress regulation in daily life.
These studies demonstrate how the CAN-STRESS dataset’s real-world applicability, ecologically valid design, and rich labeling have facilitated groundbreaking research across a range of disciplines. Its contributions underscore the value of field-based datasets in advancing our understanding of cannabis consumption, stress, and physiological monitoring.
VI Conclusion
In this paper, we introduced CAN-STRESS, a multimodal dataset designed to advance research into the physiological and behavioral effects of cannabis use in real-world settings [^1] https://zenodo.org/records/14842061. By integrating self-reported data on key daily activities, such as cannabis use, exercise, sleep patterns, and subjective stress ratings, with high-resolution physiological signals collected via a wearable wristband, CAN-STRESS provides a unique opportunity to examine the relationship between subjective experiences and objective measurements. Addressing the limitations of lab-based studies, CAN-STRESS offers an ecologically valid foundation for exploring stress regulation, activity recognition, and other health-related research domains. We encourage researchers to utilize CAN-STRESS to drive advancements in behavioral science, wearable computing, and medical diagnostics, paving the way for new insights into cannabis use and its broader implications.